# Preservation of inequalities under Hadamard products

### v1

@@ Abstract.
Wagner (1992) proved that the Hadamard product of two Pólya frequency sequences that are interpolated by polynomials is again a Pólya frequency sequence. We study the preservation under Hadamard products of related properties of significance in combinatorics. In particular, we show that ultra log-concavity, $\gamma$-positivity, and interlacing symmetric decompositions are preserved. Furthermore, we disprove a conjecture by Fischer and Kubitzke (2014) concerning the real-rootedness of Hadamard powers.
@@ 1. Introduction
Formal power series are ubiquitous in enumerative combinatorics and related areas, where the Hadamard product of generating functions often describes fundamental operations of the structures they are enumerating. A prime example is the Ehrhart series of a lattice polytope whose coefficients are given by the number of lattice points inside the dilates of the polytope by nonnegative integer factors; see, e.g., [BR15]. The Hadamard product of two Ehrhart series then equals the Ehrhart series of the Cartesian product of the corresponding lattice polytopes. More generally, if $a$ and $b$ are the Hilbert series of standard graded algebras then $a*b$ is the Hilbert series of their Segre product; see, e.g.,  [FK14].
@@
In the present paper we prove results in the same spirit of Theorem 1.1, with a particular focus on properties related to unimodality and log-concavity of the numerator polynomial defined further below. In particular, the recent development of the theory of Lorentzian polynomials by Brändén and Huh [BH20] and Anari et al. [ALOGV24] provides powerful new tools to approach log-concavity questions whose full potential still needs to be explored. We contribute to this development by applying them to Hadamard products.
@@
The following inequalities are central for the present work. A polynomial with nonnegative coefficients $\smash{a(x)=\sum_{j=0}^{d}a_{j}x^{j}}$ is called unimodal if there exists an index $k$ for which $a_{0}\leq\cdots\leq a_{k}\geq\cdots\geq a_{d}$. It is called log-concave whenever $\smash{a_{j}^{2}\geq a_{j-1}a_{j+1}}$ for every $11\leq j\leq d-$ 1. We say that $a(x)$ has no internal zeros if there are no indices $0\leq i<j<k\leq d$ such that $a_{j}=0$ and $a_{i}a_{k}\neq 0$. From the definitions it follows that any log-concave polynomial with no internal zeros is unimodal. For $m\geq d$, we say that $a(x)$ is ultra log-concave of order $m$ if the sequence
@@
is log-concave. Following Liggett [Lig97], we denote the set of ultra log-concave polynomials of order $m$ with no internal zeros by $\operatorname{ULC}(m)$. Notice that the Newton inequalities (see, e.g. [Sta89, Theorem 2]) imply that any polynomial $a(x)$ having only nonpositive real zeros is contained in $\operatorname{ULC}(\deg a)$, while the converse is easily seen to be false. In summary, if $a(x)$ is any polynomial with nonnegative coefficients, then real-rootedness (i.e., having only real zeros) implies ultra log-concavity of order $\deg a$ with no internal zeros, which in turn implies unimodality. Wagner’s Theorem 1.1 shows preservation of the most restrictive of these properties, while from [FH24, Example 3.5] it is known that unimodality is in general not preserved under the Hadamard product. In the present work, we shall investigate the middle grounds.
Throughout this paper, we shall always assume that polynomials of the form $\mathscr{W}(p)$ have nonnegative coefficients, a property that is also preserved by the Hadamard product as shown by Fischer and Juhnke-Kubitzke [FK14, Proposition 5.1(i)]); we recover this result in Lemma 4.3 below.
@@
For the proof we rely crucially on the theory of Lorentzian polynomials developed in [BH20]; we briefly review the necessary definitions and properties in Section 2. Moreover, in a sense, our result refines Wagner’s theorem, because our proof simultaneously also yields an independent proof of Theorem 1.1. We also show that the Hadamard product preserves the property of having no internal zeros; see Proposition 4.4 below. Within Ehrhart theory, this ties into results by Hofscheier, Katthän and Nill [HKN18] who showed that spanning lattice polytopes, which are closed under Cartesian products, have $h^{\ast}$-polynomials with no internal zeros. Further, in Example 2.7, we provide polynomials that show that the preservation of ultra log-concavity may fail if one removes the assumption on no internal zeros.
@@
For a polynomial $p$ of degree $d$ with $\mathcal{I}_{s}\mathscr{W}(p)=\mathscr{W}(p)$, $s\leq d$, let us define its defect, denoted $\operatorname{def}(p)$, as $d-s$. In Ehrhart theoretic terms, if $E_{P}$ is the Ehrhart polynomial of a Gorenstein polytope then $\operatorname{def}(E_{P})$ equals the degree of the lattice polytope $P$ plus $1$. As our second main result we show that $\gamma$-positivity is preserved under Hadamard products, provided both factors have the same defect, thus offering a further analog of Theorem 1.1 for $\gamma$-positivity.
@@
Our third main contribution concerns the preservation of properties for symmetric decompositions under Hadamard products. For every polynomial $h$ of degree at most $d$, there exist unique symmetric polynomials $a$ and $b$ such that $h(x)=a(x)+x\,b(x)$, $a=\mathcal{I}_{d}(a)$, and $b=\mathcal{I}_{d-1}(b)$. The pair $(a,b)$ is called the symmetric decomposition of $h$. In Ehrhart theory, symmetric decompositions were successfully used by Stapledon [Sta09] to give combinatorial proofs for linear inequalities satisfied by the coefficients of $h^{\ast}$-polynomials due to Stanley [Sta91] and Hibi [Hib90]. We say that $h$ has a nonnegative symmetric decomposition if both $a$ and $b$ have only nonnegative coefficients. Similarly, we say that the symmetric decomposition is $\gamma$-positive if both $a$ and $b$ are $\gamma$-positive. We prove the following.
@@
We will provide a counterexample to this conjecture, and even to various stronger versions of it.
@@
In fact, as we shall see, there exists an Ehrhart polynomial $p$ such that $\mathscr{W}(p^{n})$ is not log-concave for any $n\geq 1$.
@@ 2. Ultra log-concavity
where $\mathbb{N}$ denotes the set of non-negative integers, and $x^{\alpha}:=\prod_{i=1}^{n}x_{i}^{\alpha_{i}}$. The support of $p$ is defined by
@@
A polynomial $p\in\mathbb{R}[x_{1},\ldots,x_{n}]$ is said to be stable if $\operatorname{Im}(x_{i})>0$ for all $1\leq i\leq n$ implies that $p(x_{1},\ldots,x_{n})\neq 0$, i.e., $p$ is non-vanishing when all the variables take values in the upper open complex half-plane. The set of homogeneous stable polynomials of degree $d$ having non-negative coefficients is denoted by $\mathrm{S}_{n}^{d}\subseteq\mathrm{H}_{n}^{d}$.
@@
In the bivariate case, the sets of Lorentzian polynomials $\mathrm{L}_{2}^{d}\subseteq\mathbb{R}[x,y]$ and homogeneous stable polynomials with non-negative coefficients $\mathrm{S}_{2}^{d}\subseteq\mathbb{R}[x,y]$ can be described in a succinct way as follows. For every $p\in\mathrm{H}_{2}^{d}$ we have
[…]
@@
If $T$ is a homogeneous linear map that preserves stable polynomials and polynomials with non-negative coefficients, then $T$ preserves Lorentzian polynomials.
We will also make use of the fact that the following fundamental operations preserve the property of being stable or Lorentzian.
@@
Let $p$ and $q$ be polynomials with non-negative coefficients.
@@
    [BH20, Theorem 2.10] If $p(x_{1},\ldots,x_{m})$ is Lorentzian (resp. stable) and $A\in\mathbb{R}^{m\times n}$ is a matrix with non-negative entries, then $p(A\cdot(x_{1},\ldots,x_{n})^{T})$ is Lorentzian (resp. stable).
@@
    [BH20, Corollary 2.32] If $p$ and $q$ are Lorentzian (resp. stable) then so is their product $pq$.
@@
Thus, Theorem 1.2 (resp. Theorem 1.1) is equivalent to proving that if $p$ and $q$ are Lorentzian (resp. stable with non-negative coefficients), $p\bullet q$ is Lorentzian (resp. stable with non-negative coefficients). In order to do that, we begin by first explicitly describing the action of the bilinear map $(p,q)\mapsto p\bullet q$ on monomials.
@@
Specializing $m=x+a-k$, $n=x+b-\ell$, $p=a$, $q=b$, and changing the variable of the summation by $i=a+j-\ell$ we obtain
@@
which agrees with the right hand side of (2.2) after substituting $i=b+k-j$. ∎
@@
By equation (2.1), the preservation of the Lorentzian property (resp. stability) under the bilinear map $\bullet$-- ∙ - is equivalent to preservation of ultra log-concavity (resp. real-rootedness) under the Hadamard product. Thus, Theorems 1.2 and Theorem 1.1 follow immediately from Theorem 2.6. ∎
@@
The assumption on not having internal zeros in Theorem 1.2 is essential. Consider the polynomials
@@ 3.1. Hadamard products and $\gamma$-positivity
Let $p$ be a polynomial of degree $d$, denote $h=\mathscr{W}(p)$, and assume that $\deg h\leq s$. Then, the following are equivalent.
@@
Now assume that $p$ is a polynomial of degree $d$, and that $\mathscr{W}(p)$ is symmetric with center of symmetry $s/2$ for $s\leq\deg p$, i.e., $\mathcal{I}_{s}\mathscr{W}(p)=\mathscr{W}(p)$. We define the defect of $p$ as the quantity $\deg(p)-s$. We will denote this number by $\operatorname{def}(p)$. Notice that if $c=\operatorname{def}(p)$, then Lemma 3.1 implies
@@
We are now able to proof our second main result, namely that symmetry and $\gamma$-positivity is preserved under Hadamard products, given that both factors have the same defect.
@@
Let us denote $d=\deg p$ and $e=\deg q$. By Proposition 3.1, since $\mathscr{W}(p)$ and $\mathscr{W}(q)$ are symmetric, we have
@@
where $c=\operatorname{def}(f)=\operatorname{def}(q)$. Thus, $\mathscr{W}(pq)$ is again symmetric by Proposition 3.1, and evidently $\operatorname{def}(pq)=c$. Regarding $\gamma$-positivity, by bilinearity it suffices to prove the claim for
$\mathscr{W}(p)$ $=x^{i}(x+1)^{s-2i}$ for $\text{for $i\leq s$},$
$\mathscr{W}(q)$ $=x^{j}(x+1)^{\ell-2j}$ for $\text{for $j\leq\ell$},$
where $s=d-c$ and $\ell=e-c$. In that case, both $\mathscr{W}(p)$ and $\mathscr{W}(q)$ are real-rooted, and thus by Theorem 1.1, $\mathscr{W}(pq)$ is real-rooted. Since $\mathscr{W}(pq)$ is symmetric, we have that $\mathscr{W}(pq)$ is $\gamma$-positive. ∎
@@ 3.2. Ultra log-concavity of $\gamma$-polynomials
We make an interleave in our discussion about Hadamard products. A current trend in matroid theory is that of proving various log-concavity results for polynomials that arise from Hodge theory. Two families of polynomials that are conjectured to be real-rooted are the Kazhdan–Lusztig and the $Z$-polynomial of a matroid [GPY17]. The log-concavity for these two polynomials remains open even in very simple cases, though it is known that the $Z$-polynomial is palindromic and $\gamma$-positive [FMSV24]. In recent work of Xie, Wu, and Zhang [WXZ24] it was proved that for uniform matroids both the $Z$-polynomial and its associated $\gamma$-polynomial are ultra log-concave. We take the opportunity to show that there is a general implication between the ultra log-concavity of the $\gamma$-polynomial any symmetric polynomial and the ultra log-concavity of the polynomial itself (Theorem 3.3).

### v2

@@ Abstract.
Wagner (1992) proved that the Hadamard product of two Pólya frequency sequences that are interpolated by polynomials is again a Pólya frequency sequence. We study whether related combinatorial properties are preserved under Hadamard products. In particular, we show that ultra log-concavity, $\gamma$-positivity, and interlacing symmetric decompositions are preserved. Furthermore, we disprove a conjecture by Fischer and Kubitzke (2014) concerning the real-rootedness of Hadamard powers.
@@ 1. Introduction
Formal power series are ubiquitous in enumerative combinatorics and related areas, where the Hadamard product of generating functions often describes fundamental operations of the structures they are enumerating. A prime example is the Ehrhart series of a lattice polytope whose coefficients are given by the number of lattice points inside the dilates of the polytope by nonnegative integer factors; see, e.g., [BR15]. The Hadamard product of two Ehrhart series then equals the Ehrhart series of the Cartesian product of the corresponding lattice polytopes. More generally, if $a$ and $b$ are the Hilbert series of standard graded algebras then aba*b is the Hilbert series of their Segre product; see, e.g.,  [FK14].
@@
In the present paper we prove results in the same spirit of Theorem 1.1, with a particular focus on properties related to unimodality and log-concavity of the numerator polynomial defined further below. The recent development of the theory of Lorentzian polynomials by Brändén and Huh [BH20] and Anari et al. [ALOGV24] provides powerful new tools to approach log-concavity questions whose full potential still needs to be explored. We contribute to this development by applying these tools to Hadamard products.
@@
The following inequalities are central for the present work. A polynomial with nonnegative coefficients $\smash{a(x)=\sum_{j=0}^{d}a_{j}x^{j}}$ is called unimodal if there exists an index $k$ for which $a_{0}\leq\cdots\leq a_{k}\geq\cdots\geq a_{d}$. It is called log-concave whenever $\smash{a_{j}^{2}\geq a_{j-1}a_{j+1}}$ for every $1\leq j\leq d-1$. We say that $a(x)$ has no internal zeros if there are no indices $0\leq i<j<k\leq d$ such that $a_{j}=0$ and $a_{i}a_{k}\neq 0$. From the definitions it follows that any log-concave polynomial with no internal zeros is unimodal. For $m\geq d$, we say that $a(x)$ is ultra log-concave of order $m$ if the sequence
@@
is log-concave. Following Liggett [Lig97], we denote the set of ultra log-concave polynomials of order $m$ with no internal zeros by $\operatorname{ULC}(m)$. Notice that the Newton inequalities (see, e.g., [Sta89, Theorem 2]) imply that any polynomial $a(x)$ having only nonpositive real zeros is contained in $\operatorname{ULC}(\deg a)$, while the converse is easily seen to be false. In summary, if $a(x)$ is any polynomial with nonnegative coefficients, then real-rootedness (i.e., having only real zeros) implies ultra log-concavity of order $\deg a$ with no internal zeros, which in turn implies unimodality. Wagner’s Theorem 1.1 shows preservation of the most restrictive of these properties, while from [FH24, Example 3.5] it is known that unimodality is in general not preserved under the Hadamard product. In the present work, we shall investigate the middle grounds.
Throughout this paper, we shall always assume that polynomials of the form $\mathscr{W}(p)$ have nonnegative coefficients, a property that is also preserved by the Hadamard product as shown by Fischer and Juhnke-Kubitzke [FK14, Proposition 5.1(i)]; we recover this result in Lemma 4.3 below.
@@
For the proof we rely crucially on the theory of Lorentzian polynomials developed in [BH20]; we briefly review the necessary definitions and properties in Section 2. Moreover, in a sense, our result refines Wagner’s theorem, because our proof simultaneously also yields an independent proof of Theorem 1.1. We also show that the Hadamard product preserves the property of having no internal zeros; see Proposition 4.4 below. Within Ehrhart theory, this is related to results by Hofscheier, Katthän and Nill [HKN18] who showed that spanning lattice polytopes, which are closed under Cartesian products, have $h^{\ast}$-polynomials with no internal zeros. Further, in Example 2.7, we provide polynomials that show that the preservation of ultra log-concavity may fail if one removes the assumption on no internal zeros.
@@
For a polynomial $p$ of degree $d$ with $\mathcal{I}_{s}\mathscr{W}(p)=\mathscr{W}(p)$, let us define its defect, denoted $\operatorname{def}(p)$, as dsd-s. In Ehrhart theoretic terms, if $E_{P}$ is the Ehrhart polynomial of a Gorenstein polytope then $\operatorname{def}(E_{P})$ equals the degree of the lattice polytope $P$ plus $1$. As our second main result we show that $\gamma$-positivity is preserved under Hadamard products, provided both factors have the same defect, thus offering a further analog of Theorem 1.1 for $\gamma$-positivity.
@@
Our third main contribution concerns the preservation of properties for symmetric decompositions under Hadamard products. For every polynomial $h$ of degree at most $d$, there exist unique symmetric polynomials $a$ and $b$ such that h(x)=a(x)+xb(x)h(x)=a(x)+x\,b(x), $a=\mathcal{I}_{d}(a)$, and $b=\mathcal{I}_{d-1}(b)$. The pair $(a,b)$ is called the symmetric decomposition of $h$. In Ehrhart theory, symmetric decompositions were successfully used by Stapledon [Sta09] to give combinatorial proofs for linear inequalities satisfied by the coefficients of $h^{\ast}$-polynomials due to Stanley [Sta91] and Hibi [Hib90]. We say that $h$ has a nonnegative symmetric decomposition if both $a$ and $b$ have only nonnegative coefficients. Similarly, we say that the symmetric decomposition is $\gamma$-positive if both $a$ and $b$ are $\gamma$-positive. We prove the following.
@@
We will provide a counterexample to this conjecture, and even to several weaker variants of the conjecture.
@@
In fact, there exists an Ehrhart polynomial $p$ such that $\mathscr{W}(p^{n})$ is not log-concave for any $n\geq 1$.
@@ 2. Ultra log-concavity
where $\mathbb{N}$ denotes the set of nonnegative integers, and $x^{\alpha}:=\prod_{i=1}^{n}x_{i}^{\alpha_{i}}$. The support of $p$ is defined by
@@
A polynomial $p\in\mathbb{R}[x_{1},\ldots,x_{n}]$ is said to be stable if $\operatorname{Im}(x_{i})>0$ for all $1\leq i\leq n$ implies that $p(x_{1},\ldots,x_{n})\neq 0$, i.e., $p$ is non-vanishing when all the variables take values in the upper open complex half-plane. The set of homogeneous stable polynomials of degree $d$ having nonnegative coefficients is denoted by $\mathrm{S}_{n}^{d}\subseteq\mathrm{H}_{n}^{d}$.
@@
In the bivariate case, the sets of Lorentzian polynomials $\mathrm{L}_{2}^{d}\subseteq\mathbb{R}[x,y]$ and homogeneous stable polynomials with nonnegative coefficients $\mathrm{S}_{2}^{d}\subseteq\mathbb{R}[x,y]$ can be described in a succinct way as follows. For every $p\in\mathrm{H}_{2}^{d}$ we have
$p\in\mathrm{S}_{2}^{d}\iff$ $p(x,1)\text{ is a real-rooted polynomial},$
@@
If $T$ is a homogeneous linear map that preserves stable polynomials and polynomials with nonnegative coefficients, then $T$ preserves Lorentzian polynomials.
We will use the following fundamental operations, which preserve stability and the Lorentzian property
@@
Let $p$ and $q$ be polynomials with nonnegative coefficients.
@@
    [BH20, Theorem 2.10] If $p(x_{1},\ldots,x_{m})$ is Lorentzian (resp. stable) and $A\in\mathbb{R}^{m\times n}$ is a matrix with nonnegative entries, then $p(A\cdot(x_{1},\ldots,x_{n})^{T})$ is Lorentzian (resp. stable).
@@
    [BH20, Corollary 2.32] If $p$ and $q$ are Lorentzian (resp. stable) then so is their product pqpq.
@@
Thus, Theorem 1.2 (resp. Theorem 1.1) is equivalent to proving that if $p$ and $q$ are Lorentzian (resp. stable with nonnegative coefficients), $p\bullet q$ is Lorentzian (resp. stable with nonnegative coefficients). In order to do that, we begin by first explicitly describing the action of the bilinear map $(p,q)\mapsto p\bullet q$ on monomials.
@@
Specializing m=x+akm=x+a-k, $n=x+b$-$\ell$, $p=a$, $q=b$, and changing the variable of the summation by $i=a+j$-$\ell$ we obtain
@@
which agrees with the right hand side of (2.2) after substituting i=b+kji=b+k-j. ∎
@@
By equation (2.1), the preservation of the Lorentzian property (resp. stability) under the bilinear map -$\bullet$- is equivalent to preservation of ultra log-concavity (resp. real-rootedness) under the Hadamard product. Thus, Theorems 1.2 and Theorem 1.1 follow immediately from Theorem 2.6. ∎
@@
The condition that the polynomials lack internal zeros is essential in Theorem 1.2. Consider the polynomials
@@ 3.1. Hadamard products and $\gamma$-positivity
Let $p$ be a polynomial of degree $d$, denote $h=\mathscr{W}(p)$, and assume that $\deg h\leq s\leq d$. Then, the following are equivalent.
@@
Remark 3.2.
There is a ‘dual’ version of the previous lemma for the case $d\leq s\leq 2d$. In that scenario, the conditions (i) and (ii) are equivalent under the assumption that $h_{0}=h_{1}=\ldots=h_{s-d-1}=0$. The proof is analogous.
Recall that the vanishing of the coefficients $h_{i}$ for $s<i\leq d$ is equivalent to $p(-1)=p(-2)=\cdots=p(-(d-s))=0$. Dually, if $s>d$, the vanishing of coefficients $h_{i}$ for i<sdi<s-d is equivalent to $p(0)=p(1)=\cdots=p(s-d-1)=0$.
Now assume that $p$ is a polynomial of degree $d$, and that $\mathscr{W}(p)$ is symmetric with center of symmetry $s/2$, i.e., $\mathcal{I}_{s}\mathscr{W}(p)=\mathscr{W}(p)$. We define the defect of $p$ as the quantity $\deg(p)-s$. We will denote this number by $\operatorname{def}(p)$.
Observe that the defect is a negative number if $d<s\leq 2d$. For example, if $p(x)=\tfrac{1}{6}x(x-1)(x+1)$, one obtains $h(x):=\mathscr{W}(p)(x)=x^{2}$ and, in the notation above, one has $\mathcal{I}_{4}h(x)=h(x)$, so that $d=3$, $s=4$ and $\operatorname{def}(p)=3-4=-1$.
Notice that if $c=\operatorname{def}(p)$, then Lemma 3.1 (or Remark 3.2) implies
@@
We are now able to prove our second main result, namely that symmetry and $\gamma$-positivity is preserved under Hadamard products, given that both factors have the same defect.
@@
Let us denote $d=\deg p$ and $e=\deg q$. By Lemma 3.1, since $\mathscr{W}(p)$ and $\mathscr{W}(q)$ are symmetric, we have
@@
where $c=\operatorname{def}(p)=\operatorname{def}(q)$. Further, pq(i)=p(i)q(i)=0pq(-i)=p(-i)q(-i)=0 for $1\leq i\leq c$ (or $pq(i)=p(i)q(i)=0$ for $0\leq i\leq-c-1$ in the dual setting when $s>d$ and $c$ is negative; see Remark 3.2.) Thus, $\mathscr{W}(pq)$ is again symmetric by Lemma 3.1, and evidently $\operatorname{def}(pq)=c$. Regarding $\gamma$-positivity, by bilinearity it suffices to prove the claim for
$\mathscr{W}(p)$ $=x^{i}(x+1)^{s-2i}$ $\text{for $i\leq s/2$},$
$\mathscr{W}(q)$ $=x^{j}(x+1)^{\ell-2j}$ $\text{for $j\leq\ell/2$},$
where s=dcs=d-c and $\ell=e-c$. In that case, both $\mathscr{W}(p)$ and $\mathscr{W}(q)$ are real-rooted, and thus by Theorem 1.1, $\mathscr{W}(pq)$ is real-rooted. Since $\mathscr{W}(pq)$ is symmetric, we have that $\mathscr{W}(pq)$ is $\gamma$-positive. ∎
@@ 3.2. Ultra log-concavity of $\gamma$-polynomials
We make a brief digression in our discussion. A current trend in matroid theory is that of proving various log-concavity results for polynomials that arise from Hodge theory. Two families of polynomials that are conjectured to be real-rooted are the Kazhdan–Lusztig and the $Z$-polynomial of a matroid [GPY17]. The log-concavity for these two polynomials remains open even in very simple cases, though it is known that the $Z$-polynomial is palindromic and $\gamma$-positive [FMSV24]. In recent work of Xie, Wu, and Zhang [WXZ24] it was proved that for uniform matroids both the $Z$-polynomial and its associated $\gamma$-polynomial are ultra log-concave. We take the opportunity to show that there is a general implication between the ultra log-concavity of the $\gamma$-polynomial of a symmetric polynomial and the ultra log-concavity of the polynomial itself (Theorem 3.4).